OpenStudy (anonymous):

12,14,18,26,38,62,.......... what is next term in this series?

OpenStudy (mathmate):

Hint: List the first difference between the terms, then the second difference (i.e. difference between the first differences). You will see something interesting.

OpenStudy (anonymous):

multiply n add 1*2 =2 so 12+2 =14 1*4 = 4 so 14+4=18 n so on so 6*2 =12 so 62+12 =74

OpenStudy (anonymous):

multiply by 2 so when you get to 64 you will see that you multiplied 12 by 2 to get 24 so 24*2=48, 62+48=110(next term)

OpenStudy (mathmate):

I would go with aajugdar's solution.